A bus travels 10.7 miles. if the bus driver averages 45mph, how long (in minutes) will it take to travel this distance?
1
Expert's answer
2018-09-12T09:49:08-0400
We can find the time that the bus needs to travel this distance from the formula:
t=s/v=(10.7 miles)/(45 mph)=0.24 h∙(60 min)/(1 h)=14 min.